if joe has $2.35 and has only dimes and nickels with a total of 26 coins, how many dimes and how many nickels does joe have?
Answer by checkley79(3341)

235=10D+5N
D+N=26
D=26-N
235=10(26-N)+5N
235=260-10N+5N
235=260-5N
5N=260-235
5N=25
N=25/5
N=5 NICKLES.
D=26-5
D=21 DIMES.
PROOF:
235=21*10+5*5
235=210+25
235=235